# María Isabel Herrero1, Gabriela Jeronimo2,3,4 and Juan Sabia3,4

### v1

@@ 4.2 Algorithmic deformation to classify points
where $N_{\mathcal{B}}=\sum_{i=0}^{n}|\mathcal{B}_{i}|$, $\mathcal{Q}_{\mathcal{B}}=\max\{||b||:b\in\mathcal{B}_{i},i=0,\dots n\}$,
@@
(see Lemma 6 above), and according to [23, Proposition 6.1 and Section 6.2], the deformation step can be achieved within complexity $O((n^{2}N_{\mathcal{B}}\log(\mathcal{Q_{\mathcal{B}}})+n^{\Omega+1})M(D_{\mathcal{B}})M(E_{\mathcal{B}}))$ over the base field $\mathbb{Q}(s)$. Now, if $H_{\mathcal{B}}$ is an upper bound for the degrees of the numerators and the denominators of the coefficients of the polynomials in the geometric resolution we are looking for, the complexity over $\mathbb{Q}$ of the deformation step is bounded by
@@
  • $\mathcal{Q}_{\mathcal{B}}\leq\delta$ (since for every $a=(a_{1},\dots,a_{n})\in(\mathbb{Z}_{\geq 0})^{n}$, $||a||\leq a_{1}+\cdots+a_{n}$),
@@
$O(n^{4}\delta\log(\delta)NM(\delta\mathcal{D})M(\delta\mathcal{E})^{2}).$
$\square$
@@
Let $f_{1},\dots,f_{n}\in\mathbb{Q}[x_{1},\dots,x_{n}]$ be polynomials supported on $\mathcal{A}_{1},\dots,\mathcal{A}_{n}$ in $(\mathbb{Z}_{\geq 0})^{n}$ that define an algebraic variety $V\subset\mathbb{A}^{n}$, and let $\mathcal{P}\subset V$ be a finite set given by a geometric resolution $M,v_{1},\dots,v_{n}$ in $\mathbb{Q}[x_{0}]$. DiscardLowerDim is a probabilistic algorithm that, for a fixed $k$, $10\leq k\leq n-$ 1, computes a geometric resolution $\overline{M},\overline{v}_{1},\dots,\overline{v}_{n}$ in $\mathbb{Q}[x_{0}]$ (with $\overline{M}$ a factor of $M$) for a subset $\mathcal{Q}\subset\mathcal{P}$ such that $\left(V_{k}\setminus\big{(}\bigcup_{j>k}V_{j}\big{)}\right)\cap\mathcal{P}\subset\mathcal{Q}\subset\left(\bigcup_{j\geq k}V_{j}\right)\cap\mathcal{P}$. The complexity of the algorithm is of order $O(n^{4}\delta\log(\delta)N\,M(\delta\mathcal{D})M(\delta\mathcal{E})^{2})$, where $\delta$ is an upper bound for the degrees of the polynomials $M,f_{1},\dots,f_{n}$; $N=\sum_{l=1}^{n}|\mathcal{A}_{l}|$, and $\mathcal{D}$ and $\mathcal{E}$ are the parameters defined in Equations (3) and (4), respectively.

### v2

@@ 4.2 Algorithmic deformation to classify points
[…]
@@
(see Lemma 6 above), and according to [23, Proposition 6.1 and Section 6.2], the deformation stage can be achieved within complexity $O((n^{2}N_{\mathcal{B}}\log(\mathcal{Q_{\mathcal{B}}})+n^{\Omega+1})M(D_{\mathcal{B}})M(E_{\mathcal{B}}))$ over the base field $\mathbb{Q}(s)$. Now, if $H_{\mathcal{B}}$ is an upper bound for the degrees of the numerators and the denominators of the coefficients of the polynomials in the geometric resolution we are looking for, the complexity over $\mathbb{Q}$ of the deformation stage is bounded by
@@
  • $\mathcal{Q}_{\mathcal{B}}\leq\delta$ (since for every $(a_{1},\dots,a_{n})\in(\mathbb{Z}_{\geq 0})^{n}, aa$$||\leq a_{1}+\cdots+a_{n}$),
@@
$O(n^{4}\delta\log(\delta)NM(\delta\mathcal{D})M(\delta\mathcal{E})^{2}).$ $\square$
@@
Let $f_{1},\dots,f_{n}\in\mathbb{Q}[x_{1},\dots,x_{n}]$ be polynomials supported on $\mathcal{A}_{1},\dots,\mathcal{A}_{n}$ in $(\mathbb{Z}_{\geq 0})^{n}$ that define an algebraic variety $V\subset\mathbb{A}^{n}$, and let $\mathcal{P}\subset V$ be a finite set given by a geometric resolution $M,v_{1},\dots,v_{n}$ in $\mathbb{Q}[x_{0}]$. DiscardLowerDim is a probabilistic algorithm that, for a fixed $k$, $0\leq k\leq n-1$, computes a geometric resolution $\overline{M},\overline{v}_{1},\dots,\overline{v}_{n}$ in $\mathbb{Q}[x_{0}]$ (with $\overline{M}$ a factor of $M$) for a subset $\mathcal{Q}\subset\mathcal{P}$ such that $\left(V_{k}\setminus\big(\bigcup_{j>k}V_{j}\big)\right)\cap\mathcal{P}\subset\mathcal{Q}\subset\left(\bigcup_{j\geq k}V_{j}\right)\cap\mathcal{P}$. The complexity of the algorithm is of order $O(n^{4}\delta\log(\delta)N\,M(\delta\mathcal{D})M(\delta\mathcal{E})^{2})$, where $\delta$ is an upper bound for the degrees of the polynomials $M,f_{1},\dots,f_{n}$; $N=\sum_{l=1}^{n}|\mathcal{A}_{l}|$, and $\mathcal{D}$ and $\mathcal{E}$ are the parameters defined in Equations (3) and (4), respectively.